OpenStudy (help!!!!):

Find the geometric mean in the sequence 6,_,96

OpenStudy (help!!!!):

\[A. \pm24 b. 4 c.\pm16 d.51\]

OpenStudy (help!!!!):

A B C or D

OpenStudy (help!!!!):

explain the steps please

jimthompson5910 (jim_thompson5910):

first term: 6 second term: 6r third term: 6r^2

jimthompson5910 (jim_thompson5910):

the third term is 96, so 6r^2 = 96 solve for r

OpenStudy (help!!!!):

how did you find this out??

jimthompson5910 (jim_thompson5910):

just based off what a geometric sequence is and how it works

jimthompson5910 (jim_thompson5910):

each term is found by multiplying the previous term by the common ratio r

jimthompson5910 (jim_thompson5910):

btw r is not the final answer

jimthompson5910 (jim_thompson5910):

r will help you get the answer though

OpenStudy (help!!!!):

kk

OpenStudy (help!!!!):

4!!!

OpenStudy (help!!!!):

really thanks for the help

jimthompson5910 (jim_thompson5910):

the answer isn't 4

jimthompson5910 (jim_thompson5910):

but it will help you get the answer

OpenStudy (help!!!!):

oh

OpenStudy (help!!!!):

24

jimthompson5910 (jim_thompson5910):

closer

OpenStudy (help!!!!):

wait i dont understand the equation you gave me

OpenStudy (help!!!!):

but what i did was i multiplied both of them and i square rooted them

jimthompson5910 (jim_thompson5910):

r = 4 or r = -4 6r = 6*4 = 24 or 6r = 6(-4) = -24

jimthompson5910 (jim_thompson5910):

so the term that goes in the blank is 24 or -24

OpenStudy (help!!!!):

kk

jimthompson5910 (jim_thompson5910):

so that's why choice A is the answer

OpenStudy (help!!!!):

okey know I really understand lol

jimthompson5910 (jim_thompson5910):

that's great